OpenStudy (anonymous):

I really do not understand this at all. n!/(n - k)! = n(n - 1)...(n - k + 1) true or false

OpenStudy (anonymous):

It is false @torontoXO

OpenStudy (anonymous):

Can you explain @dpasingh

OpenStudy (anonymous):

Ok...

OpenStudy (anonymous):

\[\frac{n!}{(n - k)!} =\frac{n (n-1)!}{(n - k)(n-k-1)!} =\frac{n (n-1)(n-2)!}{(n - k)(n-k-1)(n-k-2)!} \] thus from the above result you can easily conclude that n!/(n - k)! = n(n - 1)...(n - k + 1) is false. @torontoXO

OpenStudy (anonymous):

Ohh ok. Understandable. Thanks bro much appreciation